OpenStudy (anonymous):

Which point is the vertex of the parabola? a: (-2, -3) b: (2, 3) c: (-2, 3) d: (2, -3)

OpenStudy (solomonzelman):

It all depends on the parabola....

OpenStudy (anonymous):

thats all they gave me

OpenStudy (anonymous):

well they gave me this too

OpenStudy (solomonzelman):

what this?

OpenStudy (anonymous):

y = x2 + 4x + 1

OpenStudy (solomonzelman):

oh, that is much better!

OpenStudy (solomonzelman):

y=x²+4x+1 you have to complete the square in this case.

OpenStudy (solomonzelman):

Let me ask you this, if you had \(\large\color{black}{ \displaystyle y=x^2+4x+\text{___} }\) then what number would you want in place of a blank to make this trinomial/expression a perfect square?

OpenStudy (anonymous):

1

OpenStudy (solomonzelman):

no, you wouldn't want a "1"..... you need this expression to factor into two equal parts.

OpenStudy (solomonzelman):

well (x+2)(x+2) would give you the first term x² and second term 4x and a third term 4, is this right or not?

OpenStudy (anonymous):

right